OpenStudy (anonymous):

Solving limits with trig identities. It has been more than 6 years since my precalc class and I need help breaking down the identities and algebraic parts of these limits. My question is lim x -> 0 of sinpix / tansqrt3x. I have gotten as far as sinpix * cossqrt3x/sinsqrt3x . If anyone has the time to show me step by step that would be so awesome. I couldn't thank you enough!

OpenStudy (anonymous):

\[\lim _{x \rightarrow 0} \sin(\Pi*x)/\tan(\sqrt(3*x))=0\]

OpenStudy (anonymous):

Thank you Ackhat, I appreciate your response. So, that wouldn't be considered indeterminate form?

OpenStudy (anonymous):

i used Lophital's rule